For Monday morning's staff meeting, Jim bought 2 bags of bagels and 3 packages of cream cheese and paid $12.00 (excluding sales tax). For Friday's meeting, he bought 4 bags of bagels and 2 packages of cream cheese and paid $14.00 (again, excluding sales tax). How much do bags of bagels and packages of cream cheese cost?

Answer by fractalier(6550)

Let x be the cost of bagels and y be the cost of cream cheese.
So the first day, he spent
2x + 3y = 12
and the second day
4x + 2y = 14
Now we solve this system by doubling the first equation and subtracting the second one from it...
2(2x + 3y = 12)
-(4x + 2y = 14)
and get
4x + 6y = 24
-(4x + 2y = 14)
and get
4y = 10
y = 2.50
Now we plug this in to either original equation (I'll use the first one) and get
2x + 3(2.50) = 12
2x + 7.50 = 12
2x = 4.50
x = 2.25
so bagels cost $2.25 and cream cheese is $2.50